OpenStudy (anonymous):

Use Part 1 of the FTC to find the derivative of each of the following functions: (a.) g(x)=integral (from 1 to x) ln(t)dt (b) F(x)=integral (from x to 10) tanthetadtheta (c)h(x)=integral (from 0 to x^2) (1+sqrt(r))dr

OpenStudy (anonymous):

ln(x) put a negative in front to get - integral from 10 to x of tan theta d theta, to get -tanx substitute u =x^2 for the last one, gtg

OpenStudy (amistre64):

FTC\[\int_{a(x)}^{b(x)}f(t)~dt=F(b(x))-F(a(x))\] taking the derivative we get \[D_x[F(b(x))-F(a(x))]=f(b(x))b'(x)-f(a(x))a'(x)\]

OpenStudy (amistre64):

or to cut the clutter \[\frac d{dx}~\int^{b}_{a}f(t)~dt=f(b)b'-f(a)a'\]